OpenStudy (anonymous):

The length of a rectangle is 2 meters more than it's width.If the area is 63 meters,what are the length and width?

OpenStudy (anonymous):

l = w+ 2 lw = 63 l(l - 2) = 63 l2 - 2l = 63 l = -7, 9 But it can't be negative, so disregard the -7. Since l = 9, w = 7

OpenStudy (anonymous):

Thanks for your help! I have the same answer just wanted to double check it!

OpenStudy (anonymous):

Welcome!